how much water could a pot hold in gallons if the pot is 10 inches tall and 12 inches in diameter
Answer by mananth(16946)

Volume of cylinder= pi*r^2*h
r=radius=6
h= height=10
Volume = pi*6^2*10
volume = 1131 in^3
231 cu.in = 1 gallon
1131/231=4.9 gallons